write in slope-intercept form the equation of the line that passes through the given point and has the given slope (1/3,-5), m=8
Answer by haileytucki(390)

((1)/(3),-5),m=8
Find the value of b using the formula for the equation of a line.
y=mx+b
Substitute the value of m into the equation.
y=(8)*x+b
Substitute the value of x into the equation.
y=(8)*((1)/(3))+b
Substitute the value of y into the equation.
(-5)=(8)*((1)/(3))+b
Since b is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
(8)*((1)/(3))+b=(-5)
Multiply (8) by ((1)/(3)) to get (8)((1)/(3)).
(8)((1)/(3))+b=(-5)
Remove the parentheses around the expression -5.
(8)((1)/(3))+b=-5
Multiply 8 by (1)/(3) to get (8)/(3).
((8)/(3))+b=-5
Reorder the polynomial (8)/(3)+b alphabetically from left to right, starting with the highest order term.
b+(8)/(3)=-5
Find the value of b.
b=-(23)/(3)
Now that the values of m(slope) and b(y-intercept) are known, substitute them into y=mx+b to find the equation of the line.
y=8x-(23)/(3)